OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle. cosine of pi divided by five times cosine of pi divided by seven plus sine of pi divided by five times sine of pi divided by seven.

OpenStudy (anonymous):

ignore that, ill post a screenshot

OpenStudy (mathmale):

Where have you seen cos a cos b + sin a sin b before? What does it represent?

jimthompson5910 (jim_thompson5910):

Hint: look on page 2 of this PDF http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf look at the "Sum and Difference Formulas" section

OpenStudy (mathmale):

Can you think of an equivalent expression for cos a cos b + sin a sin b?

OpenStudy (mathmale):

Think: trig identities!

OpenStudy (anonymous):

sin (a+b)?

OpenStudy (mathmale):

Hello, @plorb! You're on the way, but have a ways to go yet. Double check the accuracy of your answer.

OpenStudy (anonymous):

oh is it cos (a+b)?

OpenStudy (mathmale):

Double check that also. Also look at cos (a-b). Which one is it?

OpenStudy (mathmale):

cos (a+b) = ? cos (a-b) + ?

OpenStudy (anonymous):

cos a+b = cos a + cos b + sin a + sin b

OpenStudy (mathmale):

Afraid not. cos (a+b)=cos a cos b - sin a sin b. cos (a-b) + ?

OpenStudy (anonymous):

im so lost ,sorry

OpenStudy (mathmale):

sorry. cos (a-b) = ?

jimthompson5910 (jim_thompson5910):

@plohrb do you see the "Sum and Difference Formulas" section of that PDF I posted?

OpenStudy (anonymous):

yeah, im following the sum and difference formula but mathmale said it wasnt correct so im confused :(

OpenStudy (mathmale):

I'm very sorry. Look up on the 'Net: Sum and difference formulas for sine and cosine The point I've been trying to convey is that cos (a-b) = cos a cos b + sin a sin b.

OpenStudy (mathmale):

It's important that you make a list of this and other trig identities for frequent reference.

OpenStudy (anonymous):

i see

jimthompson5910 (jim_thompson5910):

ok hopefully you see this on that PDF \[\Large \cos(\alpha \pm \beta) = \cos(\alpha)\cos(\beta) \mp \sin(\alpha)\sin(\beta)\]

OpenStudy (anonymous):

yes thats the one ive been looking at

OpenStudy (mathmale):

Now apply that formula, that is, the formula for cos (a-b), to the given expression (you poosted that as an image).

OpenStudy (mathmale):

posted, not poosted. ;)

jimthompson5910 (jim_thompson5910):

the symbol \(\Large \pm\) means "plus or minus". Notice how the "plus" is on top the symbol \(\Large \mp\) means "minus or plus". The "minus" is now on top the reason why these two symbols are used is that if you have a plus on the left side, then you'll have a minus on the right side. Or vice versa

OpenStudy (anonymous):

im still very lost, what does the formula apply to the problem ?

jimthompson5910 (jim_thompson5910):

see attached

OpenStudy (anonymous):

okay?

jimthompson5910 (jim_thompson5910):

do you see how we have a pattern of "cos * cos ... sin * sin" ?

OpenStudy (anonymous):

yes i see that

jimthompson5910 (jim_thompson5910):

so that's why we're using that identity: it matches up

jimthompson5910 (jim_thompson5910):

there is a plus on the right side, so you have to use a minus on the left side

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now we substitute? or what

jimthompson5910 (jim_thompson5910):

what are \(\Large \alpha\) and \(\Large \beta\) going to be?

OpenStudy (anonymous):

5 and 7?

jimthompson5910 (jim_thompson5910):

btw \(\Large \alpha\) = greek letter alpha \(\Large \beta\) = greek letter beta

jimthompson5910 (jim_thompson5910):

not 5 and 7, but you're on the right track in a way

OpenStudy (anonymous):

pi/5 and pi/7?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

you replace all of the argument

OpenStudy (anonymous):

yay

OpenStudy (mathmale):

@plohrb : Why not write the given equation, and then, below it, write the identity that Jim has been discussing. Then it should be easier to identify the values of his alpha and beta.

OpenStudy (mathmale):

I'm so glad you have already identified your alpha and beta correctly.

OpenStudy (anonymous):

i just stated the alpha and beta right?

OpenStudy (mathmale):

Yes, right.

OpenStudy (mathmale):

So, now you want to write cos (alpha - beta), substituting the values you've identified.

OpenStudy (anonymous):

cos (pi/5 - pi/7)

OpenStudy (mathmale):

If alpha is Pi/5 and beta is Pi/7, what is the difference? subtract the latter from the former.

OpenStudy (mathmale):

Yes. You could either leave your answer like that or use the LCD to combine the two angles into one.

OpenStudy (anonymous):

wait so thats the simpe answer? or do i have to the sin part as well

OpenStudy (mathmale):

Try this: \[combined.\angle=\pi(1/5 - 1/7)\]

jimthompson5910 (jim_thompson5910):

yep you got it @plohrb \[\Large \cos\left(\alpha - \beta\right) = \cos\left(\alpha\right)\cos\left(\beta\right) + \sin\left(\alpha\right)\sin\left(\beta\right)\] \[\Large \cos\left({\color{red}{\alpha}} - {\color{blue}{\beta}}\right) = \cos\left({\color{red}{\alpha}}\right)\cos\left({\color{blue}{\beta}}\right) + \sin\left({\color{red}{\alpha}}\right)\sin\left({\color{blue}{\beta}}\right)\] \[\Large \cos\left({\color{red}{\frac{\pi}{5}}} - {\color{blue}{\frac{\pi}{7}}}\right) = \cos\left({\color{red}{\frac{\pi}{5}}}\right)\cos\left({\color{blue}{\frac{\pi}{7}}}\right) + \sin\left({\color{red}{\frac{\pi}{5}}}\right)\sin\left({\color{blue}{\frac{\pi}{7}}}\right)\] now you just simplify the expression \(\Large \frac{\pi}{5}-\frac{\pi}{7}\)

OpenStudy (mathmale):

No, you don't need the sine function here.

OpenStudy (anonymous):

ah, so the answer is just the pi/5 - pi/7

OpenStudy (mathmale):

Technically, yes. But you are asked to express your result in terms of JUST ONE angle. Thus, please combine Pi/5 - Pi/7. Alternatively, evaluate this?

OpenStudy (mathmale):

\[\pi(\frac{ 1 }{ 5 }-\frac{ 1 }{ 7})\]

OpenStudy (mathmale):

Hint: the LCD here is 5*7, or 35.

OpenStudy (anonymous):

oh um pi/35?

jimthompson5910 (jim_thompson5910):

\(\Large \frac{\pi}{5}-\frac{\pi}{7}\) is NOT equal to \(\Large \frac{\pi}{35}\). Try again

OpenStudy (mathmale):

Please show all your steps. Your numerator is incorrect.

OpenStudy (anonymous):

2pi / 35?

OpenStudy (mathmale):

Rewrite both Pi/5 and Pi/7 so that they have the LCD, which is 35. Note that you must also modify the numerators. Great job!!!1

OpenStudy (mathmale):

So, what is your final answer? Please go back and look at the question before responding.

OpenStudy (anonymous):

cos = 2pi/35.

OpenStudy (mathmale):

"Write the following in terms of the sine, or cosine, or tangent, of a single angle. Take out that " = " symbol and write what's left. 2pi/35 is the "argument" of the cosine function.

OpenStudy (mathmale):

The given expression, involving the sine and cosine, can be expressed as the ________ (name of trig function) of a single angle, and that angle is ___________.

OpenStudy (anonymous):

its just cos 2pi/35 then?

OpenStudy (mathmale):

Perfect. Congrats!

jimthompson5910 (jim_thompson5910):

I'd use parenthesis and say `cos(2pi/35)`

jimthompson5910 (jim_thompson5910):

but yes you have the answer

OpenStudy (mathmale):

You will see these trig identies again and again.

OpenStudy (mathmale):

Worth writing them down and reviewing them regularly so that you know them immediately when you need them.

OpenStudy (anonymous):

thanks to both of you! have a great evening :)

OpenStudy (mathmale):

My great pleasure. And thanks so much to you, Jim, for your thoughtful input.